# Novelty and Impact of Economics Papers

### v1

@@ E.3 Robustness of Disruption Analysis
Novelty Archetype Disruption Rate Std 95% CI lower 95% CI upper
Consolidating 0.1184 0.0034 0.1118 0.1250
Outlying 0.2245 0.0032 0.2181 0.2308
Trendy 0.1053 0.0024 0.1006 0.1100
Trailblazing 0.2021 0.0037 0.1948 0.2093
all 0.1630 0.0017 0.1597 0.1663
@@
Novelty Archetype Disruption Rate Std 95% CI lower 95% CI upper
Consolidating 0.1038 0.0065 0.0912 0.1165
Outlying 0.2012 0.0073 0.1869 0.2154
Trendy 0.1062 0.0051 0.0962 0.1162
Trailblazing 0.1694 0.0052 0.1591 0.1797
all 0.1451 0.0029 0.1393 0.1509
@@
Novelty Archetype Disruption Rate Std 95% CI lower 95% CI upper
Consolidating 0.1327 0.0136 0.1060 0.1593
Outlying 0.2271 0.0122 0.2031 0.2510
Trendy 0.1219 0.0071 0.1080 0.1359
Trailblazing 0.1819 0.0082 0.1658 0.1979
all 0.1638 0.0048 0.1545 0.1732
@@ E.4 Robustness of the Advantage of Trailblazing Archetype
Novelty Archetype Dual Impact Rate Std. Dev. 95% CI Lower 95% CI Upper
Consolidating 0.0188 0.0012 0.0165 0.0211
Outlying 0.0396 0.0016 0.0365 0.0428
Trendy 0.0309 0.0015 0.0280 0.0339
Trailblazing 0.0550 0.0017 0.0516 0.0584
all 0.0359 0.0007 0.0345 0.0374
@@
Novelty Archetype Dual Impact Rate Std. Dev. 95% CI Lower 95% CI Upper
Consolidating 0.0086 0.0009 0.0068 0.0104
Outlying 0.0167 0.0009 0.0149 0.0185
Trendy 0.0144 0.0009 0.0127 0.0161
Trailblazing 0.0245 0.0012 0.0222 0.0267
all 0.0159 0.0005 0.0150 0.0169
@@ E.5 Robustness to Alternative Novelty Threshold
Novelty Archetype High-Cite Rate Std 95% CI lower 95% CI upper
Consolidating 0.2136 0.0018 0.2100 0.2172
Outlying 0.2188 0.0038 0.2114 0.2262
Trendy 0.3401 0.0039 0.3325 0.3477
Trailblazing 0.3756 0.0097 0.3565 0.3946
all 0.2477 0.0001 0.2475 0.2479
@@
Novelty Archetype Disruption Rate Std 95% CI lower 95% CI upper
Consolidating 0.1442 0.0024 0.1395 0.1488
Outlying 0.2508 0.0040 0.2430 0.2586
Trendy 0.1150 0.0031 0.1090 0.1210
Trailblazing 0.2075 0.0091 0.1897 0.2254
all 0.1630 0.0017 0.1597 0.1663

### v3

@@ E.3 Robustness of Disruption Analysis
Neighborhood Archetype Disruption Rate SE 95% CI lower 95% CI upper
Consolidating 0.1184 0.00034 0.11778 0.11910
Outlying 0.2245 0.00032 0.22385 0.22513
Trendy 0.1053 0.00024 0.10484 0.10578
Trailblazing 0.2021 0.00037 0.20135 0.20279
all 0.1630 0.00017 0.16264 0.16331
@@
Neighborhood Archetype Disruption Rate SE 95% CI lower 95% CI upper
Consolidating 0.1038 0.00065 0.10254 0.10507
Outlying 0.2012 0.00073 0.19974 0.20259
Trendy 0.1062 0.00051 0.10519 0.10719
Trailblazing 0.1694 0.00052 0.16839 0.17045
all 0.1451 0.00029 0.14453 0.14569
@@
Neighborhood Archetype Disruption Rate SE 95% CI lower 95% CI upper
Consolidating 0.1327 0.00136 0.13000 0.13533
Outlying 0.2271 0.00122 0.22467 0.22947
Trendy 0.1219 0.00071 0.12054 0.12334
Trailblazing 0.1819 0.00082 0.18028 0.18349
all 0.1638 0.00048 0.16291 0.16478
@@ E.4 Robustness of the Advantage of Trailblazing Archetype
Neighborhood Archetype Dual Impact Rate SE 95% CI Lower 95% CI Upper
Consolidating 0.0188 0.00012 0.01854 0.01900
Outlying 0.0396 0.00016 0.03932 0.03995
Trendy 0.0309 0.00015 0.03064 0.03123
Trailblazing 0.0550 0.00017 0.05466 0.05534
all 0.0359 0.00007 0.03580 0.03609
@@
Neighborhood Archetype Dual Impact Rate SE 95% CI Lower 95% CI Upper
Consolidating 0.0086 0.00009 0.00839 0.00874
Outlying 0.0167 0.00009 0.01652 0.01688
Trendy 0.0144 0.00009 0.01419 0.01453
Trailblazing 0.0245 0.00012 0.02423 0.02469
all 0.0159 0.00005 0.01585 0.01603
@@ E.5 Robustness to Alternative Novelty Threshold
Neighborhood Archetype High-Cite Rate SE 95% CI lower 95% CI upper
Consolidating 0.2136 0.00018 0.21322 0.21395
Outlying 0.2188 0.00038 0.21809 0.21957
Trendy 0.3401 0.00039 0.33932 0.34085
Trailblazing 0.3756 0.00097 0.37365 0.37746
all 0.2477 0.00001 0.24771 0.24775
@@
Neighborhood Archetype Disruption Rate SE 95% CI lower 95% CI upper
Consolidating 0.1442 0.00024 0.14370 0.14463
Outlying 0.2508 0.00040 0.25004 0.25161
Trendy 0.1150 0.00031 0.11441 0.11561
Trailblazing 0.2075 0.00091 0.20576 0.20933
all 0.1630 0.00017 0.16264 0.16331